OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (−2, 6) and has a slope of −1? x + y = −4 x + y = 4 x − y = −4 x − y = 4

ganeshie8 (ganeshie8):

write it in slope-intercept form first

ganeshie8 (ganeshie8):

then, convert it to standard

ganeshie8 (ganeshie8):

scratch that, write it in point-slope form first

ganeshie8 (ganeshie8):

point : (−2, 6) slope : −1

ganeshie8 (ganeshie8):

point slope form :- y-y1 = m(x-x1)

OpenStudy (anonymous):

y - 6 = -1(x- 2)

ganeshie8 (ganeshie8):

care ful of signs, its -2

ganeshie8 (ganeshie8):

y-6 = -1(x--2) y-6 = -1(x+2) y-6 = -x - 2 x+y = 4

OpenStudy (anonymous):

alight i got it

OpenStudy (anonymous):

thanks man